# Spatial-Network Treatment Effects: A Continuous Functional Approach

### v3

@@ 3.5.3 Results: Binary Treatment Timing
Figure 2 presents event study plots averaging across 1,000 Monte Carlo replications for each simulation scenario. Panel A in Figure 2 shows results when the true treatment effect structure has no spillovers. All estimators perform similarly, tracking the true effect closely. The pre-treatment coefficients are centered at zero, confirming that all methods satisfy parallel trends when SUTVA holds. The no-spillover PDE and continuous GPS slightly outperform TWFE due to their use of continuous treatment intensity, but differences are modest. Panel B in Figure 2 shows results when the true treatment effect structure includes spillovers. The patterns diverge dramatically: TWFE underestimates the true effect by approximately 25% at all post-treatment horizons. This bias arises because TWFE attributes spillovers to the “control” group, attenuating the treatment-control contrast. Continuous GPS underestimates by approximately 20%. Although GPS uses continuous treatment intensity, it conditions only on own treatment, missing the spillover channel. Units with low own-treatment but high spillover exposure are misclassified. No-spillover PDE (our framework with $\nu_{s}=\nu_{n}=0$ imposed) underestimates by approximately 30%. Imposing the wrong restriction forces the model to attribute spillover-driven outcomes to direct effects, biasing the decay parameter $\kappa$ and overall effect magnitude. Full PDE correctly tracks the true effect throughout the post-treatment period. By estimating $(\nu_{s},\nu_{n},\lambda)$ rather than imposing them to zero, the full model captures both direct effects and spillovers.
Table 2 presents results for binary treatment timing (treatment switches on at $t=5$). When the true model has no spillovers, all estimators perform well with low bias and correct coverage. The Full PDE correctly detects that $\nu_{s}\approx\nu_{n}\approx 0$. When spillovers are present, conventional methods exhibit substantial bias (12–38%) and severe undercoverage (52–84%). Only the Full PDE maintains low bias and correct coverage across all scenarios.
Table 2: Monte Carlo Results: Binary Treatment Timing
True Model Estimator Bias RMSE Coverage
No spillovers TWFE 0.02 0.08 0.94
Continuous GPS 0.01 0.07 0.95
No-spillover PDE 0.00 0.06 0.96
Full PDE 0.01 0.07 0.95
Spatial only TWFE 0.18-0.18 0.21 0.72
Continuous GPS 0.14-0.14 0.17 0.78
No-spillover PDE 0.22-0.22 0.25 0.68
Full PDE 0.02 0.09 0.94
Network only TWFE 0.12-0.12 0.15 0.81
Continuous GPS 0.09-0.09 0.12 0.84
No-spillover PDE 0.15-0.15 0.18 0.76
Full PDE 0.01 0.08 0.95
Full model TWFE 0.31-0.31 0.34 0.58
Continuous GPS 0.25-0.25 0.28 0.65
No-spillover PDE 0.38-0.38 0.41 0.52
Full PDE 0.02 0.10 0.94
  • Notes: 1,000 replications. Bias and RMSE for long-run effect. Coverage is 95% CI rate. Full PDE uses spatial-network HAC; others use clustered SEs.
3.5.4 Results: Continuous Treatment Intensity
The framework’s key advantage is accommodating continuous treatment intensity. We modify the simulation to draw treatment from a log-normal distribution with intensity varying by location and market position.
Table 3: Monte Carlo Results: Continuous Treatment Intensity
True Model Estimator Parameter Bias RMSE Coverage
No spillovers TWFE (binarized) ATE 0.08 0.12 0.89
Continuous GPS $\mu(1)$ 0.01-0.01 0.08 0.95
No-spillover PDE $\beta_{s}$ 0.01 0.05 0.96
Full PDE $\beta_{s}$ 0.01 0.06 0.95
Full model TWFE (binarized) ATE 0.22 0.26 0.68
Continuous GPS $\mu(1)$ 0.12 0.16 0.76
No-spillover PDE $\beta_{s}$ 0.25 0.28 0.62
Full PDE $\beta_{s}$ 0.02 0.07 0.94
Full PDE $\beta_{n}$ 0.02-0.02 0.08 0.95
Full PDE $\lambda$ 0.02-0.02 0.06 0.94
  • Notes: 1,000 replications. TWFE binarizes treatment at the median. Full PDE recovers all structural parameters.
[…]
3.5.5 Heterogeneous Marginal Effects
The true marginal effect under the full model is:
$\frac{\partial\tau}{\partial S}=\beta_{s}+\lambda N$ (74)
This varies with network exposure NN—units with high network exposure experience larger marginal effects. Conventional methods estimate constant marginal effects, missing this heterogeneity.
00.20.20.40.40.60.60.80.8111.21.21.41.40.40.40.60.60.80.8111.21.21.41.41.61.6Network Exposure (NN)Marginal Effect ($\partial\tau/\partial S$)True: 0.8+0.4N0.8+0.4NTWFE: constantGPS: constantFull PDE: $0.82+0.38N$
Figure 3: Heterogeneous Marginal Effects by Network Exposure
Notes: True marginal effect increases with network exposure. Conventional methods estimate constant effects, missing heterogeneity. Full PDE correctly recovers the slope.
Figure 3 illustrates the heterogeneity. The true marginal effect (black) increases linearly with network exposure. TWFE and GPS estimate constant effects (horizontal lines), missing the fundamental heterogeneity. The Full PDE correctly recovers both the intercept and slope.
3.5.6 Key Findings from Monte Carlo
The Monte Carlo evidence establishes three findings:
Nested structure: When spillovers are absent, the Full PDE correctly detects this and produces estimates equivalent to conventional methods. The framework does not spuriously find spillovers where none exist.
Bias reduction: When spillovers are present, conventional methods exhibit 25–38% bias while the Full PDE maintains bias under 3%. The bias in conventional methods arises from attributing spillover-driven outcomes to direct effects or to the “control” group.
Correct inference: The spatial-network HAC standard errors achieve correct coverage (94–96%) while conventional standard errors achieve only 52–68% coverage when both spatial and network dependence are present.
3.6 GMM Estimation
We combine the three identification strategies within a GMM framework.
3.6.1 Moment Conditions
The moment conditions are:
$\mathbf{g}_{1}(\boldsymbol{\theta})$ $=\mathbb{E}\left[(Y_{i}-\beta_{s}S_{i}-\beta_{n}N_{i}-\lambda S_{i}N_{i}-\mathbf{X}_{i}^{\prime}\boldsymbol{\gamma})\cdot\mathbf{Z}_{i}^{\text{RD}}\right]=0$ (75)
$\mathbf{g}_{2}(\boldsymbol{\theta})$ $=\mathbb{E}\left[(Y_{i}-\beta_{s}S_{i}-\beta_{n}N_{i}-\lambda S_{i}N_{i}-\mathbf{X}_{i}^{\prime}\boldsymbol{\gamma})\cdot\mathbf{Z}_{i}^{\text{IV}}\right]=0$ (76)
$\mathbf{g}_{3}(\boldsymbol{\theta})$ $=\mathbb{E}\left[\log\hat{D}_{KL}(t)+2\lambda_{2}t-c\right]=0$ (77)
where $\mathbf{Z}_{i}^{\text{RD}}$ are spatial RD instruments, $\mathbf{Z}_{i}^{\text{IV}}$ are network instruments, and the third condition uses entropy dynamics.
The parameter vector is $\boldsymbol{\theta}=(\beta_{s},\beta_{n},\lambda,\kappa,\nu_{s},\nu_{n},\boldsymbol{\gamma})^{\prime}$.
3.6.2 Estimation
Theorem 3 (GMM Estimator Properties).
Let $\hat{\boldsymbol{\theta}}_{\text{GMM}}$ denote the two-step GMM estimator. Under standard regularity conditions:
Consistency: $\hat{\boldsymbol{\theta}}_{\text{GMM}}\xrightarrow{p}\boldsymbol{\theta}_{0}$
Asymptotic normality: $\sqrt{N}(\hat{\boldsymbol{\theta}}_{\text{GMM}}-\boldsymbol{\theta}_{0})\xrightarrow{d}N(0,\mathbf{V})$
Specification test: The Hansen $J$-statistic $J=N\hat{\mathbf{g}}^{\prime}\hat{\mathbf{W}}\hat{\mathbf{g}}\xrightarrow{d}\chi^{2}_{m-k}$ under correct specification.
The overidentification test is valuable: if the three strategies yield inconsistent estimates, the model is misspecified.
3.7 Spatial-Network HAC Inference
Valid inference requires accounting for both spatial and network dependence.
Definition 8 (Spatial-Network HAC Variance Estimator).
$\hat{\mathbf{V}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}K\left(\frac{d_{ij}^{s}}{h_{s}},\frac{d_{ij}^{n}}{h_{n}}\right)\hat{\mathbf{g}}_{i}\hat{\mathbf{g}}_{j}^{\prime}$ (78)
where $d_{ij}^{s}$ is spatial distance, $d_{ij}^{n}$ is network distance, and $K(\cdot,\cdot)$ is a product kernel.
The product kernel structure assumes spatial and network correlation decay multiplicatively:
$K(u_{s},u_{n})=k_{s}(u_{s})\cdot k_{n}(u_{n})$ (79)
Units that are both spatially proximate and network-connected receive high weight; units distant on either dimension receive low weight.
Proposition 12 (HAC Consistency).
Under mixing conditions and appropriate bandwidth growth ($h_{s},h_{n}\to\infty$, $h_{s}h_{n}/N\to 0$):
$\hat{\mathbf{V}}\xrightarrow{p}\mathbf{V}_{0}$ (80)
The Monte Carlo evidence confirms that spatial-network HAC achieves correct coverage (94–96%) when both types of dependence are present.
4 Empirical Application
We apply the framework to U.S. minimum wage policy over 2018–2023, combining QCEW wage data, BEA input-output tables, and IRS migration data.
4.1 Data
Wage data cover 3,142 counties, 274 industries, and 24 quarters (approximately 17.8 million observations). Supply chain networks are constructed from BEA input-output tables. Labor mobility networks are constructed from IRS migration data.
Treatment intensity is measured continuously:
$S_{c,t}=\max\{w^{\min}_{s(c),t}-\bar{w}_{c,t-1},0\}$ (81)
the dollar amount by which the minimum wage exceeds the lagged average wage.
4.2 Testing the No-Spillover Null
Table 4 reports tests for spillovers.
Table 4: Tests for Spillovers
Test Coefficient Std. Error p-value
Distance to border ($d$) 0.0018-0.0018 0.0004 <0.001<0.001
Network exposure ($\tilde{N}$) 0.032 0.007 <0.001<0.001
$d\times\tilde{N}$ 0.0012-0.0012 0.0003 <0.001<0.001
Joint test ($F$-statistic) 89.2 <0.001<0.001
  • Notes: Standard errors clustered by state.
All three predictions under the no-spillover null are rejected: treatment effects exhibit significant spatial gradients, network gradients, and interaction. The joint F-test decisively rejects ($p<0.001$).
4.3 Main Estimates
Table 5 presents GMM estimates.
Table 5: GMM Estimates of Spatial-Network Treatment Effects
Parameter Estimate Std. Error Interpretation
Treatment Effects
Direct effect ($\beta_{s}$) 0.041 0.007 \$/hr per \$1 MW
Network effect ($\beta_{n}$) 0.028 0.007 Supply chain spillover
Mixed effect ($\lambda$) 0.043 0.008 Interaction
Propagation Parameters
Spatial diffusion ($\nu_{s}$) 98.4 12.3 sq mi/quarter
Network diffusion ($\nu_{n}$) 0.48 0.08 market position units
Decay rate ($\kappa$) 0.28 0.03 per quarter
Derived Quantities
Spatial half-distance 35 mi $\ln(2)/\kappa_{s}$
Temporal half-life 2.5 Q $\ln(2)/\kappa$
Mutual information 0.38 nats $I(\mathbf{x};\alpha)$
Specification Test
Hansen $J$-statistic 5.2 $p=0.47$
  • Notes: Spatial-network HAC standard errors.
The estimates have economic interpretations:
Spatial diffusion $\hat{\nu}_{s}=98$ square miles per quarter implies annual mobility scale of approximately 20 miles, consistent with commuting distances.
Decay rate $\hat{\kappa}=0.28$ implies 2.5-quarter half-life, consistent with gradual labor market adjustment.
Mutual information $\hat{I}(\mathbf{x};\alpha)=0.38$ quantifies spatial clustering of supply chains.
The Hansen $J$-test passes ($p=0.47$), supporting correct specification.
4.4 Comparison with Conventional Methods
Table 6 compares methods.
Table 6: Comparison of Estimation Methods
Method Direct Spatial Network Mixed Total
TWFE 0.031 0.031
Continuous GPS 0.042 0.042
No-spillover PDE 0.044 0.044
Full PDE 0.041 0.041 0.028 0.043 0.153
[…]
Conventional methods estimate direct effects of \$0.031–0.044. The Full PDE estimates total effects of \$0.153 at the border—3.7 times larger. Conventional methods capture only 27% of total policy impact.
4.5 Decomposition
At the border, direct effects account for 27% of total; spatial spillovers 27%, network spillovers 16%, and mixed effects 30%. The mixed effect—synergistic amplification when both channels reinforce each other—accounts for nearly one-third of total impact.

### v4

@@ 3.2 Monte Carlo Evidence
Panel B shows results when the true DGP includes spatial and network spillovers with interaction. The patterns diverge dramatically. TWFE underestimates the true effect by approximately 25% at all post-treatment horizons; this bias arises because TWFE attributes spillovers to the control group, attenuating the treatment-control contrast. Continuous GPS underestimates by approximately 20%; although GPS uses continuous treatment intensity, it conditions only on own treatment, missing the spillover channel. No-spillover PDE (our framework with $\nu_{s}=\nu_{n}=0$ imposed) underestimates by approximately 30%; imposing the wrong restriction forces the model to attribute spillover-driven outcomes to direct effects, biasing the decay parameter $\kappa$ and overall effect magnitude. Only Full PDE correctly tracks the true effect throughout the post-treatment period by estimating $(\nu_{s},\nu_{n},\lambda)$ rather than imposing them to zero.
3.3 Overview of Identification Strategy
Section 3.2 identified three fundamental challenges to identifying spatial-network treatment effects: the reflection problem, spatial confounding, and network endogeneity. This subsection previews our three-pronged identification strategy and connects it to conventional methods.
3.3.0.1 Three Complementary Approaches.
We develop three identification strategies, each addressing different challenges while building on established econometric principles. The first strategy exploits spatial regression discontinuity, following Lee and Lemieux (2010) and Dube et al. (2010), by utilizing sharp treatment boundaries at state borders. This addresses spatial confounding by comparing units just on either side of borders where treatment jumps discontinuously but confounders vary smoothly. The connection to conventional methods is direct: this is standard RD extended to continuous treatment intensity, where we recover not just the discontinuity size but also the spatial decay parameter $\kappa_{s}$ from how the discontinuity attenuates with distance from the border.
The second strategy uses network instrumental variables, following Bramoullé et al. (2009), by employing predetermined network connections as instruments. Historical supply chain relationships established before treatment cannot respond to current shocks, providing exogenous variation in network exposure. This is standard two-stage least squares applied to network spillovers: the first stage predicts current network exposure using historical connections, while the second stage estimates treatment effects conditional on predicted exposure.
The third strategy leverages entropy-based moment conditions that use theoretical predictions about relative entropy decay to generate overidentifying restrictions. This approach is unique to the continuous framework and has no direct analog in discrete methods. It extends GMM by adding moment conditions from the theoretical structure of the master equation, similar to how Euler equation restrictions provide overidentification in consumption models, but applied to the distributional dynamics of treatment effects.
3.3.0.2 Advantages of Multiple Strategies.
Combining three identification approaches provides several advantages over relying on a single method. Triangulation operates when all three strategies yield similar estimates, strengthening confidence in the results, while disagreement signals misspecification and directs attention to which assumptions may be violated. Efficiency gains arise because GMM combining all moment conditions is more efficient than any single estimator, with the Hansen $J$-test using the overidentification to test model specification.
Different sources of variation are exploited by each approach: RD uses border discontinuities, IV uses historical network connections, and entropy uses distributional predictions. Each exploits different variation in the data, reducing dependence on any single identification assumption. Robustness is enhanced because some applications may lack clean RD designs (no sharp borders) or adequate instruments (networks too stable), while having multiple strategies allows implementation even when one approach fails.
3.3.0.3 Comparison with Conventional Methods.
Table 5 compares our approach to conventional treatment effect methods. The table reveals a fundamental distinction: conventional methods cannot identify spillover parameters because they maintain SUTVA. Spatial extensions relax SUTVA for spatial spillovers but ignore networks. Network extensions relax SUTVA for network spillovers but ignore space. Our framework relaxes SUTVA for both dimensions simultaneously by combining the three complementary strategies.
Table 5: Identification Strategies: Conventional vs. Our Framework
Method Variation Exploited Key Assumption Can ID Spillovers?
Conventional Methods
TWFE Within-unit over time Parallel trends No
DiD Treatment timing Parallel trends No
GPS Continuous dose Unconfoundedness No
Matching Covariate balance Ignorability No
Spatial Extensions
Spatial DiD Timing + distance Parallel + smooth Spatial only
Border RD Sharp boundaries Continuity at border Spatial only
Spatial HAC Inference only
Network Extensions
Network IV Historical network Exclusion + relevance Network only
Peer effects Group structure Partial interference Network only
Our Framework
Spatial RD Border discontinuities Continuity Yes (both)
Network IV Historical connections Exclusion + relevance Yes (both)
Entropy GMM Distributional dynamics Theoretical prediction Yes (both)
Combined GMM All three sources All three jointly Yes (both)
  • Notes: Conventional methods cannot identify spillover parameters. Spatial extensions identify spatial spillovers only; network extensions identify network spillovers only. Our framework identifies both simultaneously.
3.4 Relationship Between GMM and Conventional Panel Estimators
The GMM framework may appear complex relative to standard panel regression. This subsection clarifies the relationship and shows that conventional panel estimators are special cases of our GMM approach.
3.4.0.1 Panel Regression as Restricted GMM.
Standard panel regression with two-way fixed effects:
$Y_{it}=\beta S_{it}+\alpha_{i}+\gamma_{t}+\varepsilon_{it}$ (103)
can be understood as GMM with moment conditions:
$\mathbb{E}[(Y_{it}-\beta S_{it}-\alpha_{i}-\gamma_{t})\cdot Z_{it}]=0$ (104)
where the instrument set $Z_{it}$ includes $S_{it}$ and indicator variables for units and time periods.
Our GMM framework extends this specification in five ways. First, we add spatial exposure variables defined as $\tilde{S}_{it}^{s}=\sum_{j}w_{ij}^{s}S_{jt}$ where $w_{ij}^{s}$ is a spatial weight function. Second, we add network exposure defined as $\tilde{S}_{it}^{n}=\sum_{j}G_{ij}S_{jt}$ using network adjacency $G_{ij}$. Third, we add interaction exposure $\tilde{S}_{it}^{\lambda}=\sum_{j}w_{ij}^{s}G_{ij}S_{jt}$ to capture spatial-network complementarities. Fourth, we use additional instruments from spatial RD (discontinuities at borders) and network IV (historical connections). Fifth, we add entropy-based overidentifying restrictions that test the consistency of parameter estimates with theoretical predictions about distributional dynamics.
When $\nu_{s}=\nu_{n}=0$, the spatial and network exposure terms drop out, additional instruments become irrelevant, and entropy restrictions are satisfied trivially. In this case, our GMM moment conditions reduce exactly to those of standard panel regression in equation (103). The GMM framework therefore nests conventional panel methods as a testable special case.
3.4.0.2 Efficiency Gains.
Proposition 3.2 (Efficiency of Combined GMM).
Let $\hat{\beta}_{\text{RD}}$, $\hat{\beta}_{\text{IV}}$, and $\hat{\beta}_{\text{GMM}}$ denote estimates from spatial RD alone, network IV alone, and combined GMM respectively. Under correct specification:
$\text{Var}[\hat{\beta}_{\text{GMM}}]\leq\min\{\text{Var}[\hat{\beta}_{\text{RD}}],\text{Var}[\hat{\beta}_{\text{IV}}]\}$ (105)
with equality only if moment conditions are redundant.
This follows from the standard GMM efficiency result: using more moment conditions weakly improves efficiency. The combined GMM uses moment conditions from RD, IV, and entropy jointly, hence is weakly more efficient than any individual approach. The efficiency gain can be substantial when individual instruments are weak but jointly strong, a common situation in spatial-network settings where border discontinuities or historical network connections alone provide limited variation but together identify parameters precisely.
3.4.0.3 Decision Framework.
For applied researchers, the choice between methods depends on the research question and data availability. Standard panel regression as in equation (103) is appropriate when primary interest focuses on direct treatment effects only, when spillovers are theoretically implausible or empirical tests fail to reject $H_{0}:\nu_{s}=\nu_{n}=0$, when data lack spatial coordinates or network structure necessary for constructing exposure variables, or when the audience is unfamiliar with GMM methodology and simpler presentation is valued.
Our GMM framework becomes appropriate when spillovers are theoretically plausible and empirical tests reject the no-spillover null, when policy evaluation requires decomposing total effects into direct versus spillover components, when data have rich spatial and network structure enabling construction of exposure variables, when multiple identification strategies are available (RD designs, valid instruments, entropy calculations), or when efficiency gains from combining approaches are important for detecting economically meaningful effects.
For most applications in spatial economics, international trade, and public finance where spillovers are likely important, the GMM framework is appropriate. For applications where SUTVA is reasonable, such as randomized experiments with distant units or policies affecting isolated markets, conventional methods suffice and should be preferred for their simplicity.
3.4.0.4 Computational Requirements.
Table 6 compares computational requirements between standard panel regression and our GMM framework. Panel regression using standard software packages (reghdfe in Stata, fixest in R, PanelOLS in Python) requires only outcome, treatment, and covariate data, runs in 2–5 seconds for typical datasets with 3,000 units and 20 time periods, needs 10–20 lines of code, and estimates only direct treatment effects with clustered standard errors but no overidentification tests or policy counterfactuals.
Table 6: Computational Requirements: Panel Regression vs. GMM
Aspect Panel Regression Our GMM Framework
Software reghdfe/fixest/PanelOLS Custom implementation
Required inputs $(Y,S,X)$ $(Y,S,X,\text{coords},G)$
Time (N=3000, T=20) 2–5 seconds 2–3 minutes
Lines of code 10–20 100–200
Estimates Direct effect only Direct + spillovers + interaction
Standard errors Clustered Spatial-network HAC
Overidentification No Hansen $J$-test
Policy counterfactuals No Yes (via Feynman-Kac)
  • Notes: Computational time on standard laptop. Panel regression is faster but GMM provides richer output. For practitioners needing only point estimates, discrete panel regression (Section 4) provides fast approximation.
Our GMM framework requires custom implementation code, needs additional inputs for spatial coordinates and network adjacency matrix, takes 2–3 minutes for the same dataset size, and requires 100–200 lines of code for complete implementation. However, it estimates direct effects plus spatial spillovers, network spillovers, and their interaction, computes spatial-network HAC standard errors accounting for both dependence sources, provides Hansen $J$-test for overidentification, and enables policy counterfactuals via the Feynman-Kac representation without re-estimation.
The GMM framework requires more computational effort but provides substantially more information about treatment propagation. For researchers who need the additional structure for spillover decomposition or policy counterfactuals, the extra implementation time is worthwhile. For practitioners seeking only point estimates of direct effects, the discrete panel regression approximation in Section 4 provides a middle ground: standard tools, fast computation, but estimates that capture spillovers if present.
3.5 Summary
This section has developed identification and estimation methods for spatial-network treatment effects. Before proceeding to the empirical application, we summarize how our approach extends conventional econometric practice.
3.5.0.1 What We Retain from Conventional Methods.
Our identification strategies build on three established principles. First, regression discontinuity exploits sharp boundaries where treatment jumps while other factors vary smoothly, a principle dating to Thistlethwaite and Campbell (1960) and formalized by Hahn et al. (2001). Second, instrumental variables use predetermined variables to address endogeneity of exposure measures, following the tradition from Wright (1928) through modern applications. Third, generalized method of moments combines moment conditions for efficiency and provides overidentification tests, building on Hansen (1982). These are not novel contributions but are standard tools applied carefully to the spatial-network setting.
When tests fail to reject the hypothesis that spillovers are zero ($H_{0}:\nu_{s}=\nu_{n}=0$), our estimators reduce exactly to conventional approaches. Specifically, we recover the generalized propensity score method of Hirano and Imbens (2004) for continuous treatment, the doubly robust estimator of Kennedy et al. (2017) when our structural outcome model is correctly specified, and heterogeneity-robust difference-in-differences following Callaway and Sant’Anna (2021) and Sun and Abraham (2021) for staggered adoption designs. The framework nests these conventional methods as testable special cases rather than replacing them.
GMM consistency and asymptotic normality in our framework follow from standard conditions without requiring new asymptotic theory. Spatial-network HAC inference uses the same kernel-based approach as spatial HAC developed by Conley (1999), extended through product kernels to account for both spatial and network dimensions simultaneously. The theoretical foundations are conventional; the application to joint spatial-network dependence is what requires careful development.
3.5.0.2 What We Add Beyond Conventional Methods.
Four substantive extensions distinguish our framework from conventional approaches. First, simultaneous spatial and network spillovers: conventional methods address spatial OR network spillovers but not both, while our framework estimates spatial diffusion rate $\nu_{s}$, network diffusion rate $\nu_{n}$, and their interaction coefficient $\lambda$. This enables decomposing total effects into direct, spatial spillover, network spillover, and interaction components, revealing which transmission channels matter quantitatively.
Second, testable restrictions from theory: the master equation provides testable predictions about event study coefficient decay patterns (should follow exponential at rate $\kappa$), relative entropy dynamics (should decay at rate $2\lambda_{2}$), and spatial gradient functional forms (should match steady-state solutions). These restrictions have no analog in atheoretical regression models and provide additional specification tests beyond conventional diagnostics.
Third, structural interpretation and counterfactuals: parameters have clear economic meanings rather than being reduced-form regression coefficients. For example, $\nu_{s}=98$ square miles per quarter translates to a 20-mile annual labor mobility scale, $\kappa=0.28$ per quarter implies a 2.5-quarter adjustment half-life, and $\lambda=0.04$ nats indicates moderate geographic industry clustering. These structural parameters enable counterfactual policy evaluation using the Feynman-Kac representation without re-estimating the model for each policy scenario.
Fourth, aggregation-robust inference: conventional spatial models yield coefficients that vary with geographic aggregation level (ZIP code versus county versus state estimates differ substantially). Our structural parameters are aggregation-invariant, remaining constant whether estimated at fine or coarse spatial resolution. This property is essential for policy evaluation that must be robust to how geographic units are defined.
3.5.0.3 Practical Workflow.
For applied researchers, we recommend the following four-step workflow. First, start with tests for spillovers by implementing the joint test $H_{0}:\nu_{s}=\nu_{n}=\lambda=0$ from Section 3.2 using the augmented regression with spatial, network, and interaction terms. Second, if no spillovers are detected (test fails to reject), use conventional methods such as GPS for continuous treatment or DiD for binary treatment timing, as these are simpler to implement, easier to communicate to general audiences, and our framework adds no value when spillovers are absent. Third, if spillovers are detected (test rejects), use our GMM framework because it is necessary to avoid bias in direct effect estimates, enables spillover decomposition for policy analysis understanding transmission channels, and provides structural parameters for counterfactuals evaluating alternative policy designs. Fourth, for robustness checking, implement both discrete panel regression (fast, familiar to referees) and continuous GMM (efficient, structural), where agreement within 10% suggests discrete approximation suffices for point estimates, while substantial disagreement signals model misspecification requiring investigation.